OpenStudy (anonymous):

Write each answer as a power of 2. Show your work and explain your steps. a. Computer capacity is often measured in bits and bytes. A bit is the smallest unit, which is a 1 or a 0, in the computer's memory. A byte is 2^3 bits. A megabyte, (MB), is 2^20 bytes. How many bits are in a megabyte?

OpenStudy (anonymous):

i think you want \[(2^3)^20=2^{3\times 20}=2^{60}\]

OpenStudy (anonymous):

typo there \[(2^3)^{20}=2^{3\times 20}=2^{60}\]